a given mass of chlorine occupies 38cm2 at 20°c. Determine it final volume at 45°c. Pressure remains constant


1
Expert's answer
2020-11-12T09:32:38-0500

For isobaric process (constant pressure), "\frac{V_1}{T_1} = \frac{V_2}{T_2}", where the temperature is in kelvins. Hence, "V_2 = \frac{V_1 T_2}{T_1} = \frac{38 cm^3 \cdot 293.15 K}{318.15K} = 35.01 cm^3"